OpenStudy (anonymous):

if f(x) = 1 / x^3 + 2 ; what is f(y^2)

OpenStudy (amoodarya):

put x=y^2 in the function

OpenStudy (anonymous):

1 / y^(2^3) + 2?

OpenStudy (amoodarya):

yes go on

OpenStudy (anonymous):

i don't know whats next

OpenStudy (amoodarya):

1/y^6+2

OpenStudy (amoodarya):

because (y^2)^3=y^6

OpenStudy (anonymous):

yeah i get it now

OpenStudy (anonymous):

so is that the answer?